OpenStudy (anonymous):

x^2+17x/x^2-3x + x^2-10x/x^2-3x add and simplify of possible

OpenStudy (anonymous):

(2x+7)/(x-3)

OpenStudy (anonymous):

(x^2+17x+x^2-10x)/(x(x-3))=(x(2x+7))/(x(x-3))=(2x+7)/(x-3)